OpenStudy (anonymous):

find the missing term of each geometric sequence 3,__, 12

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

There are a number of ways to go about this one. Do you know the formula for the nth term of a geometric sequence?

OpenStudy (anonymous):

no well maybe \[An=a \times r^{n-1}\]

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Okay, so we have \(a_1 = 3\) and \(a_3 = 12\) if \(r\) is the common ratio, then \[a_2 = a1_*r\]\[a_3 = a_2*r = (a_1*r)*r\]

OpenStudy (whpalmer4):

\[a_3 = a_1*r^2\]\[\frac{12}{3} = r^2\]\[r=\]

OpenStudy (anonymous):

r=2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so will it be An=3 * 2^2-1???? will that be the equation we use to find the 2nd term

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

r = 2. so \[a_n = a_1*r^{n-1} = 3*2^{n-1}\]

OpenStudy (anonymous):

if we are trying to find the 2nd term then n=2 right

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

i figured it out thank you

OpenStudy (whpalmer4):

so what is the second term?

OpenStudy (anonymous):

6

OpenStudy (whpalmer4):

very good. how about the 10th term? :-)